OpenStudy (anonymous):

Simplify the expression x^2y^4 over x^-3y^7)^-2 A. y^5/x^7 B. x^3/y C. y^6/x^10 D. x^-6/y^-5 help?

OpenStudy (anonymous):

\[\frac{x^2y^4}{x^{-3}y^7}=\frac{x^{2+3}}{y^{7-4}}=\frac{x^5}{y^3}\] is a good start